a man is four timesaas old as his son. in 3 years, the father will be three times as old as his son. how old is each now?
Answer by mananth(16946)

son = x years
man = 4x
..
ages after 3 years
son x+3
man=4x+3
...
4x+3=3(x+3)
4x+3=3x+9
4x-3x=9-3
x=6 son's age
Mans age = 4*6 = 24 years
...
m.ananth@hotmail.ca